# GNN2R: Weakly-Supervised Rationale-Providing Question Answering over Knowledge Graphs

### 04 Dec 2023 v1

@@ Abstract
Knowledge Graphs, Question Answering, Graph Neural Networks, Language Models
@@ I Introduction
In recent years, knowledge graphs (KGs) have been prominently used as knowledge sources supporting multi-hop question-answering (QA) [1]. An example is given in Fig. 1, where the question “who was born in California and directed a movie starring Michael Keaton” can be answered based on the given KG with Tim Burton. This seems to be a successful application of KG-based QA. However, it is difficult to verify and understand the answer for normal users, given the complexity of navigating question-relevant parts within a large KG, not to mention the technical expertise required for such navigation, e.g., SPARQL [2] and understanding of the schema of the KG [3]. We propose to provide users with a reasoning subgraph consisting of triples from the KG that elucidate how the answer satisfies the semantic constraints specified in the question. As highlighted in the figure, the reasoning subgraph of this question consists of two paths: [(Tim Burton, birthplace, California)], which affirms that Tim Burton was born in California, and [(Tim Burton, director-1, Batman 1989),  [^1] We use the superscript -1 to denote the inverse of a relation. (Batman 1989, cast member, Michael Keaton)], which affirms that Tim Burton directed Batman 1989 starring Michael Keaton. Hence, reasoning subgraphs provide users with a rationale for the final answers. Also, in the spirit of transparent AI, intermediate information in reasoning subgraphs, e.g., the alluded movie in this example, could be useful and important.
@@ II Related Work
We introduce three types of current mainstream QA methods and discuss the potential application of large language models.
@@
Query construction-based methods [4, 5, 6, 7, 8, 9, 10] try to translate given questions into logical queries, e.g., SPARQL [2] and symbolic programs [11], that can be executed over KGs to obtain final answers. They have been studied for a long time and have achieved competitive performance. Also, by instantiating the constructed logical queries, these methods can be potentially used to generate reasoning subgraphs. However, recent methods of this type, especially case-based [6], transfer learning-based [7], and supervised learning-based [10], heavily rely on annotations of logical queries, which makes them inapplicable in low-resource scenarios with only weak supervision available, including the task of this paper, where only final answers are annotated for training.
@@
Retrieval-based methods [12, 13, 14, 15, 16, 17] compute contextual KG representations and leverage distance computations or learned neural classifiers to retrieve answers by the parametric ranking or classification of entities. For example, Sun et al. [14] propose to use a graph neural network [13] to compute contextual representations of KG entities and use these embeddings to classify entities into answers and non-answer entities. These methods are highly efficient. However, as these methods do not explicitly reason over triples, they do not offer explanations. In the first step, our model selects candidate answers from computed embedding spaces, which aligns with this category.
@@
Reasoning-based methods [18, 19, 20, 21, 22, 23] are potentially capable of extracting reasoning subgraphs from KGs. A branch [18, 20] of these methods performs explicit explorations in KGs and generates reasoning chains for given questions. For example, Qiu et al. [20] use a reinforcement learning agent that explores paths in the KG starting from topic entities mentioned in given questions. Another branch [21, 22] performs fuzzy reasoning by computing confidence scores or numerical labels for entities and triples in KGs. For example, Shi et al. [21] propose to compute stepwise scores for relations in KGs to determine how likely a triple is part of the reasoning chain of a given question. These methods are the ones most applicable to the generation of reasoning subgraphs. However, the weak supervision and efficiency issues have not been well addressed. Some methods, such as [19], rely on annotations of reasoning chains, and their mechanisms for explicit KG reasoning, such as [20], are prone to inefficiency. Furthermore, most of these methods can only provide probable reasoning chains with likelihoods that are difficult to comprehend.
@@
The application of large language models (LLMs), in KG-based QA has recently attracted wide attention [24, 25, 26]. For example, Tan et al. [25] find that LLMs have the potential to outperform traditional QA methods, which do not rely on LLMs, in a relaxed setup, where a question is deemed correctly answered if the response from LLMs mentions any valid answer. However, this study also highlights a crucial drawback of LLMs—extracting answers from their textual response of varying length is difficult, especially considering that the answers may be expressed as aliases, acronyms, synonyms, or even in a different language. Also, given the implicit nature of the knowledge in LLMs and the common hallucination issue [27], it is challenging to verify the validity and completeness of answers from LLMs. In contrast, traditional QA methods can provide explicit answers in an unambiguous and uniform manner, and our method is able to extract the rationale behind answers from KGs. Furthermore, Sun et al. [26] show that even the current state-of-the-art LLMs, including ChatGPT [^2] https://platform.openai.com/docs/api-reference and LLaMA (33B) [28], still have difficulty answering simple (one-hop) factual questions generated from templates, which is often considered a solved task with traditional QA methods. Therefore, in the era of LLMs, the contributions of this paper, alongside other methods without relying on LLMs, still hold practical significance.
@@ IV-A Step-I: GNN-based Coarse Reasoning
$\mathbf{e}^{k}=\frac{\exp\left(s^{k}_{m}-s^{\prime k}_{max}\right)\mathbf{m}^{k}+\exp\left(s^{k}_{e}-s^{\prime k}_{max}\right)\mathbf{e}^{k-1}}{\exp\left(s^{k}_{m}-s^{\prime k}_{max}\right)+\exp\left(s^{k}_{e}-s^{\prime k}_{max}\right)},$ (9)
where $s^{\prime k}_{max}=\max(s^{k}_{m},s^{k}_{e})$.
@@ IV-B Step-II: LM-based Explicit Reasoning
[…]
3:natural language expressions.
4:$\operatorname{NL}_{p}(q)\leftarrow$ an empty set
5:$\operatorname{NL}_{n}(q)\leftarrow$ an empty set
6:$nl2ans\leftarrow$ an empty dictionary
7:$nl2vote\leftarrow$ an empty dictionary
8:$A_{c}\leftarrow$ candidate answer set of $q$ computed via Step-I
9:for entity $e\in A_{c}\cup A_{q}$ do
10:     $G_{q}\leftarrow$Longunderstack[l]the set of candidate reasoning subgraphs ex-
11:tracted with $e$ using Algorithm 1
12:     for subgraph $\mathcal{G}_{c}\in G_{q}$ do
13:         $Q(\mathcal{G}_{c})\leftarrow$Longunderstack[l]convert $\mathcal{G}_{c}$ into a query by turning $e$ in-
14:to a variable
15:         $A(\mathcal{G}_{c})\leftarrow$Longunderstack[l]a set of entities that can be obtained by
16:executing $Q(\mathcal{G}_{c})$ over $\mathcal{G}$
17:         $nl\leftarrow$Longunderstack[l]the natural language expression generated
18:based on $\mathcal{G}_{c}$ using Algorithm 2
19:         if $nl\notin keys(nl2ans)$ then
20:              $nl2ans[nl]\leftarrow$ an empty set
21:         end if
22:         for entity $e^{\prime}\in A(\mathcal{G}_{c})$ do
23:              Add $e^{\prime}$ to $nl2ans[nl]$
24:         end for
25:     end for
26:end for
27:for $nl\in$ $keys(nl2ans)$ do
28:     $upvote\leftarrow\left|nl2ans[nl]\cap A_{q}\right|$
29:     $downvote\leftarrow\left|nl2ans[nl]\setminus A_{q}\right|$
30:     if $upvote>0$ then
31:         $nl2vote[nl]\leftarrow upvote-downvote$
32:     end if
33:end for
34:$max\_vote\leftarrow$Longunderstack[l]maximum vote of any $nl\in keys(nl2vote)$
35:$min\_len\leftarrow$Longunderstack[l]minimum number of entity/relation ment-
36:ions in the expressions that have $max\_vote$
37:for $nl\in$ $keys(nl2vote)$ do
[…]
39:         Add $nl$ to $\operatorname{NL}_{p}(q)$
40:     else
41:         Add $nl$ to $\operatorname{NL}_{n}(q)$
42:     end if
43:end for
@@ V-E Explainability Evaluation
Explanation Evaluation. We use ground-truth annotations of reasoning chains in PQ and PQL to further examine the quality of our generated reasoning subgraphs. A comparison with recent reasoning-based methods, including NSM, SR+NSM, and TransferNet, is reported in Table VIII. Specifically, we check if the triples within our reasoning subgraphs or the explanations of the baselines are present in the ground-truth reasoning chains and calculate precision, recall, and F1 on this triple level. It can be observed that, in general, the quality of our generated reasoning subgraphs is substantially higher than that of existing reasoning-based methods. In comparison to the best-performing baseline, i.e., TransferNet, average improvements of 52.7% (precision), 15.5% (recall), and 31.4% (F1) are achieved on the four datasets. Also, on the datasets with less training data and bigger KGs, i.e., PQL-2hop and PQL-3hop, the average improvements of GNN2R become even larger: 96.7% (precision), 29.2% (recall), and 57.8% (F1). NSM and SR+NSM are not comparable to TransferNet and our model regarding the quality of generated explanations. The main reason is that NSM and SR+NSM do not examine if their reasoning chains are semantically consistent with given questions. As for TransferNet, the major weakness is the precision of explanations. As analyzed above, TransferNet tends to return redundant relations and, consequently, triples in final explanations. This is also the reason that it performed better than GNN2R on PQ-3hop regarding recall. However, a higher recall does not necessarily signify a better quality of explanation, especially given its lower precision.

### 20 Jan 2024 v3

@@ Abstract
Knowledge Graph, Question Answering, Graph Neural Network, Language Model
@@ I Introduction
In recent years, knowledge graphs (KGs) have been prominently used to organize and provide knowledge for question answering (QA) tasks [1, 2, 3, 4, 5]. An example is given in Fig. 1, where the question “who was born in California and directed a movie starring Michael Keaton” can be answered based on the given KG with Tim Burton. This seems to be a successful application of KG-based QA. However, it is difficult to verify and understand the answer for normal users, given the complexity of navigating question-relevant parts within a large KG, not to mention the technical expertise required for such navigation, e.g., SPARQL [6] and understanding of the schema of the KG. We propose to provide users with a reasoning subgraph consisting of triples from the KG that elucidate how the answer satisfies the semantic constraints specified in the question. As highlighted in the figure, the reasoning subgraph of this question consists of two paths: [(Tim Burton, birthplace, California)], which affirms that Tim Burton was born in California, and [(Tim Burton, director-1, Batman 1989),  [^1] We use the superscript -1 to denote the inverse of a relation. (Batman 1989, cast member, Michael Keaton)], which affirms that Tim Burton directed Batman 1989 starring Michael Keaton. Hence, reasoning subgraphs provide users with a rationale for the final answers. Also, in the spirit of transparent AI, intermediate information in reasoning subgraphs, e.g., the alluded movie in this example, could be useful and important.
@@ II Related Work
In this section, we introduce three types of current mainstream KG-based QA methods and discuss the application of large language models (LLMs) in this domain.
@@
Query construction-based methods [7, 8, 9, 10] translate given questions into logical queries, e.g., SPARQL [6] and symbolic programs [11], that can be executed over KGs to obtain answers. This type of method has been extensively studied since the inception of KG-based QA. Also, by instantiating the constructed logical queries, these methods can be potentially used to generate reasoning subgraphs. However, recent SOTA methods of this type, including case-based [7], transfer learning-based [8], and supervised learning-based [10], heavily rely on existing corpora or annotations of logical queries, making them inapplicable to the task of this paper, where only weak supervision (i.e., annotations of final answers) is available for training.
@@
Retrieval-based methods [12, 13, 14, 15] directly retrieve answers from KGs via parametric ranking or classification of KG entities based on their contextual representations, e.g., KG embeddings. For example, Sun et al. [13] propose to use a graph neural network [12] to compute contextual representations of KG entities and use these embeddings to classify entities into answers and non-answer entities. These methods are usually highly efficient, as they employ end-to-end numerical computations without explicit reasoning over KGs. Consequently, they lack interpretability and cannot provide explicit answer explanations. The first step of our model selects candidate answers from computed embedding spaces, which is in line with this type of method. However, the above weaknesses are effectively mitigated by the second step.
@@
Reasoning-based methods [16, 17, 18, 19, 20, 21] can be divided into two branches. A branch [16, 18] performs explicit explorations in KGs and generates reasoning chains for given questions. For example, Qiu et al. [18] use a reinforcement learning agent that explores paths in the KG starting from topic entities mentioned in given questions. Another branch [19, 20] performs fuzzy reasoning by computing confidence scores or numerical labels for KG entities and triples. For example, Shi et al. [19] propose to compute stepwise scores for relations in KGs to determine how likely a triple is part of the reasoning chain of a given question. These methods are the ones most applicable to the generation of reasoning subgraphs. However, the weak supervision and efficiency issues have not been well addressed. Some methods, such as [17], rely on annotations of reasoning chains, and their mechanisms for explicit KG reasoning, such as [18], are prone to inefficiency. Furthermore, most of these methods can only provide probable reasoning chains with likelihoods that are difficult to comprehend.
@@
The application of LLMs in KG-based QA has recently attracted wide attention [22, 23, 24]. For example, Tan et al. [23] find that LLMs have the potential to outperform traditional QA methods that do not rely on LLMs in a relaxed setup, where a question is deemed correctly answered if the response from LLMs mentions any valid answer. However, this study also highlights a crucial drawback of LLMs—extracting answers from their textual response of varying length is difficult, especially considering that the answers may be expressed as aliases, acronyms, synonyms, or even in a different language. Also, given the implicit nature of the knowledge in LLMs and the common hallucination issue [25], it is challenging to verify the validity and completeness of answers from LLMs. In contrast, traditional QA methods can provide explicit answers in an unambiguous and uniform manner. Also, our method is able to extract the rationale behind answers from KGs. Furthermore, Sun et al. [24] show that even the current SOTA LLMs, including ChatGPT [^2] https://platform.openai.com/docs/api-reference and LLaMA (33B) [26], still have difficulty answering simple (one-hop) factual questions generated from templates, which is often considered a solved task with traditional QA methods. Therefore, in the era of LLMs, the contributions of this paper, alongside other methods without relying on LLMs, still hold practical significance.
@@ IV-A Step-I: GNN-based Coarse Reasoning
$\mathbf{e}^{k}=\frac{\exp\left(s^{k}_{m}-s^{\prime k}_{max}\right)\mathbf{m}^{k}+\exp\left(s^{k}_{e}-s^{\prime k}_{max}\right)\mathbf{W}^{k}_{t3}\mathbf{e}^{k-1}}{\exp\left(s^{k}_{m}-s^{\prime k}_{max}\right)+\exp\left(s^{k}_{e}-s^{\prime k}_{max}\right)},$ (9)
where $s^{\prime k}_{max}=\max(s^{k}_{m},s^{k}_{e})$ and $\mathbf{W}^{k}_{t3}\in\mathbb{R}^{d_{k}\times d_{k-1}}$. Also, $\mathbf{W}^{k}_{t3}$ should be an identity matrix if $d_{k}=d_{k-1}$.
@@ IV-B Step-II: LM-based Explicit Reasoning
[…]
2:answer set $A_{q}$, computed results of GNN2R Step-I.
[…]
4:$\operatorname{NL}_{n}(q)$ of natural language expressions.
5:$\operatorname{NL}_{p}(q)\leftarrow$ an empty set
6:$\operatorname{NL}_{n}(q)\leftarrow$ an empty set
7:$nl2ans\leftarrow$ an empty dictionary
8:$nl2vote\leftarrow$ an empty dictionary
9:$A_{c}\leftarrow$ candidate answer set of $q$ computed via Step-I
10:for entity $e\in A_{c}\cup A_{q}$ do
11:     $G_{q}\leftarrow$Longunderstack[l]the set of candidate reasoning subgraphs ex-
12:tracted with $e$ using Algorithm 1
13:     for subgraph $\mathcal{G}_{c}\in G_{q}$ do
14:         $Q(\mathcal{G}_{c})\leftarrow$Longunderstack[l]convert $\mathcal{G}_{c}$ into a query by turning $e$ in-
15:to a variable
16:         $A(\mathcal{G}_{c})\leftarrow$Longunderstack[l]a set of entities that can be obtained by
17:executing $Q(\mathcal{G}_{c})$ over $\mathcal{G}$
18:         $nl\leftarrow$Longunderstack[l]the natural language expression generated
19:based on $\mathcal{G}_{c}$ using Algorithm 2
20:         if $nl\notin keys(nl2ans)$ then
21:              $nl2ans[nl]\leftarrow$ an empty set
22:         end if
23:         for entity $e^{\prime}\in A(\mathcal{G}_{c})$ do
24:              Add $e^{\prime}$ to $nl2ans[nl]$
25:         end for
26:     end for
27:end for
28:for $nl\in$ $keys(nl2ans)$ do
29:     $upvote\leftarrow\left|nl2ans[nl]\cap A_{q}\right|$
30:     $downvote\leftarrow\left|nl2ans[nl]\setminus A_{q}\right|$
31:     $nl2vote[nl]\leftarrow upvote-downvote$
32:end for
33:$max\_vote\leftarrow$Longunderstack[l]maximum vote of any $nl\in keys(nl2vote)$
34:$min\_len\leftarrow$Longunderstack[l]minimum number of entity/relation ment-
35:ions in the expressions that have $max\_vote$
36:for $nl\in$ $keys(nl2vote)$ do
[…]
38:         Add $nl$ to $\operatorname{NL}_{p}(q)$
39:     else
40:         Add $nl$ to $\operatorname{NL}_{n}(q)$
41:     end if
42:end for
@@ V-E Explainability Evaluation
Explanation Evaluation. We use ground-truth annotations of reasoning chains in PQ and PQL to further examine the quality of our generated reasoning subgraphs. A comparison with recent reasoning-based methods, including NSM, SR+NSM, and TransferNet, is reported in Table VIII. Specifically, we check if the triples within our reasoning subgraphs or the explanations of the baselines are present in the ground-truth reasoning chains and calculate precision, recall, and F1 on the triple level. The following can be observed:
1) It can be observed that, except for a performance drop on PQ-3hop regarding recall, the quality of our generated reasoning subgraphs is substantially higher than that of existing reasoning-based methods. In comparison to the best-performing baseline, i.e., TransferNet, average improvements of 52.7% (precision), 15.5% (recall), and 31.4% (F1) are achieved on the four datasets. Also, on the datasets with less training data and bigger KGs, i.e., PQL-2hop and PQL-3hop, the average improvements of GNN2R become even larger: 96.7% (precision), 29.2% (recall), and 57.8% (F1). This demonstrates the strength of GNN2R in handling large search spaces with limited weak supervision.
2) NSM and SR+NSM are not comparable to TransferNet and our model regarding the quality of generated explanations. The main reason is that NSM and SR+NSM do not examine if their reasoning chains are semantically consistent with given questions. As illustrated by the above example “who is gimli’s father in the hobbit,” the reasoning chains of NSM and SR+NSM are often incorrect, despite yielding correct final answers. As for TransferNet, the major weakness is the precision of explanations. This is consistent with the above analysis that TransferNet tends to return too many relations and, consequently, redundant triples in final explanations. This contributed to a higher recall, especially on PQ-3hop. However, given its lower precision, the higher recall cannot signify a better quality of explanation.